OpenStudy (anonymous):

A ac votage source have peak valure of 10v in series of a ideal diodeand resistor,what is the V dc component accross the resistor

OpenStudy (anonymous):

the answer says 3.18v but i work out to be 5v see attachment for the original question.

OpenStudy (anonymous):

The circuit is a half wave rectifier, then: \[V_{dc}=V_p/(2\pi)\int\limits_{0}^{\pi}\sin(\phi)d \phi=V_p/\pi\]